OpenStudy (tia_sweetheart17):

if the sum of a number and nine is tripled, the result is one less than twice the number. find the number

OpenStudy (ddcamp):

You're given two sides of an equation: "the sum of a number and nine is tripled" and "one less than twice the number". Let "the number" be x. What are the two sides of the equation now?

OpenStudy (anonymous):

\[2x-3(x+9)=1\rightarrow2x-3x-27=1\rightarrow-x=28rightarrowx=-28\]

OpenStudy (anonymous):

3(x+9)=2x-1